OpenStudy (anonymous):

If f(x)=8(5x-2), find f(x)=3+4/3??? HELP!!!!!

OpenStudy (anonymous):

that thing on the right is a number is there a typo there?

OpenStudy (anonymous):

YES! f(x) =3x -4, fing f(x)=3+4/3

OpenStudy (anonymous):

wait 3x-4?

OpenStudy (anonymous):

Yes typo

OpenStudy (anonymous):

how did i guess? what is it really?

OpenStudy (anonymous):

well 3+4/3=13/3 so f(13/3)=3(13/3)-4 f(13/3)=13-4 f(13/3)=9

OpenStudy (anonymous):

Ok it is...If f(x)=3x-4, find f(x)=3+4/3

OpenStudy (anonymous):

THANK you thank you thank you!!!!!!!!